OpenStudy (anonymous):

The speed of train A is 20mph slower than the speed of train B. Train A travels 210miles in the same time it takes train B to travel 310 miles. find the speed of each train. @-------College Al…

OpenStudy (anonymous):

Is it 42mph and 62mph

OpenStudy (anonymous):

Train B: Speed = X Travels 310miles in Y amount of time 310/X = Y ------ (equation 1) Train A: Speed = X-20 Travels 210miles in Y amount of time 210/(X-20) = Y ------ (equation 2) Knowing that:\[\frac{310}{x} = \frac{210}{x-20}\]\[210x = 310x - 620\] solve for X, then plug back into the equations 1 and 2 to find each train speed